# Percolation analysis of spatio-temporal distribution of population in Seoul and Helsinki

### v1

@@ Abstract
Spatio-temporal distribution of urban population is crucial to understand the structure and dynamics of cities. Most studies, however, have focused on the microscopic structure of cities such as their few most crowded areas. In this work, we investigate the macroscopic structure of cities such as their clusters of highly populated areas. To do this, we analyze the spatial distribution of urban population and its intra-day dynamics in Seoul and Helsinki with a percolation framework. We observe that the growth patterns of the largest clusters in the real and randomly shuffled population data are significantly different, and highly populated areas during daytime are denser and form larger clusters than during nighttime. An analysis of the cluster size distributions at percolation criticality shows that their power-law exponents during daytime are lower than during nighttime, indicating that the spatial distributions of urban population during daytime and nighttime fall into different universality classes. Finally measuring the area-perimeter fractal dimension of the collection of clusters at criticality demonstrates that the fractal dimensions during daytime are higher than during nighttime, indicating that the perimeters of clusters during daytime are more rough than during nighttime. Our findings suggest that even the same city can have qualitatively different spatial distributions of population over time, and propose a way to quantitatively compare the macrostructure of cities based on population distribution data.
@@ I Introduction
Characterizing the spatio-temporal distribution of urban population is important because it not only helps understand the structure and dynamics of cities but also provides essential information on demand for traffic, housing, and infrastructure in cities. Spatial distribution of population in many cities are not uniform but quite heterogeneous [1, 2]. A key concept in urban structure, thus, is urban centers or hotspots, usually defined as the most crowded places in cities [3, 4]. Urban structure can be classified as either monocentric or polycentric based on the number of centers, and compact or dispersed based on how closely clustered centers are [5, 6].
@@
Until recently, the registered population based on censuses has been mainly used to understand the spatial distribution of urban population [1, 7, 8]. These data are highly reliable as collected by the governments, but they are updated every five to ten years and are based on a residence-based population, which limits its ability to describe the dynamics of urban structure. However, real-time or near-real-time population data collected from mobile phone users have recently become available and can be used to better understand urban structure and dynamics [9, 10, 11, 12, 13, 14]. A remarkable work proposed a parameter-free method to identify hotspots in 31 Spanish cities, showed that the number of hotspots in these cities scales sublinearly with their population, and classified these hotspots according to their lifetime [15]. Mobile phone data also allow to investigate the urban structure encoded in mobility flows of people in cities with the notion of urban centers [16, 17]. These center-based investigations focus on the microscopic structure of cities although they revealed import structural properties of cities. A comprehensive understanding of the structure and dynamics of a city requires an investigation of the macrostructure of the spatial distribution of the city’s population such as clusters of highly populated areas.
@@
Percolation can provide an effective way to investigate the macroscopic structure underlying a given spatial distribution of variables of interest from temperature and altitude to congestion and population [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. For example, a work that analyzed spatial distribution of temperature in cities with a temperature-based percolation (i.e., occupying sites in order of the highest temperature to the lowest temperature) revealed that urban heat islands are fractal despite the diversity in urban form, composition, and environment across the world [29]. Another work, using a percolation model to analyze spatial distributions of congestion in two cities, showed that these distributions belong to different universality classes of percolation depending on the commuting patterns of the cities [20].
@@
In this work, we investigate the intra-day dynamics of the macroscopic structure underlying the spatial distribution of population in Seoul and Helsinki with a percolation framework. We examine the characteristics of clusters of highly populated areas and the connectivity between these clusters as they are created by occupying areas from high to low population in this framework. We observe that the growth patterns of the largest clusters in the real-world and randomly shuffled population data are significantly different, and highly populated areas during daytime are denser and form larger clusters than during nighttime by analyzing the size of the largest clusters, the radius of gyration of occupied areas, and the correlation length of clusters in the data. We investigate the cluster size distributions at percolation critical point and observe that their power-law exponents during daytime are lower than during nighttime, indicating that the spatial distributions of urban population during daytime and nighttime belong to different universality classes. We measure the area-perimeter fractal dimension of the collection of clusters at critical point and observe that the fractal dimensions during daytime are higher than during nighttime, indicating that the perimeters of high-population clusters during daytime are more rough than during nighttime.
@@ II.1 Data
We use the Seoul Living Population Data (SLPD) [30] provided by the Seoul Metropolitan Government to analyze the spatio-temporal distribution of population in Seoul. The SLPD is based on mobile phone signal data and tells us the total number of people of a specific area in Seoul at a specific time (an hourly interval). There are 19,153 area units, called ‘jibgyegu’, in the dataset. Figure 1(a) shows the spatial distribution of Seoul’s population at 9:00 A.M. on April 1st, 2019 by the area unit.
Since the ‘jibgyegu’ area units are irregular in shape and size, we spatially aggregated the number of people recorded each hour in each area unit on a regular square grid of cell size 400m. As a result, we obtained 4043 grid cells with geographic coordinates and hourly population density. We considered weekdays in April 2019 for our analysis (20 days in total). We tested the robustness of our results with two different sizes of grid cells (250m and 500m) and obtained qualitatively similar results.
To focus on relatively high population areas in cities, we excluded grid cells whose daily population densities are lower than the daily population densities averaged over all the grid cells. Figure 1(b) shows the spatial distribution of Seoul’s population at 9:00 A.M. on April 1st, 2019 on the grid cells after the exclusion of lowly-populated cells. The number of grid cells was reduced from 4043 to 2480. The excluded lowly-populated grid cells are mostly located in mountains, rivers, and public facilities such as sewage treatment plants.
Figure 1: Spatial distribution of population density in Seoul at 9:00 AM on April 1st, 2019. The color bar represents the logarithmic values of population density in spatial units. (a) Spatial distribution by the ‘jibgyegu’ units. (b) Spatial distribution by grid cells of 400m size after the exclusion of lowly-populated cells.
@@ II.2 Population-based percolation
The procedure of our population-based percolation is as follows. We first rank all the grid cells according to their population density from the highest to the lowest value. Next, we occupy the grid cells one by one according to their rank and observe how clusters evolve with occupation probability $p$. A cluster, defined as a set of nearest neighboring occupied grid cells, represents a densely populated macroscopic area in a city. The occupation probability $p$ is defined as the number of occupied grid cells divided by the total number of grid cells.
@@ III.1 Evolution of the largest cluster and the number of clusters
Our percolation analysis aims at characterizing the macroscopic structure underlying spatial distribution of urban population by investigating the clusters of highly-populated areas in cities. First we investigate the size of the largest cluster as a function of the occupation probability $p$. Figure 2(a) and 2(b) show the size of the largest cluster at 2:00 P.M. and 2:00 A.M. on Apr. 1st, 2019 in Seoul as a function of $p$, respectively.
@@ III.2 Radius of gyration and correlation length
We considered these quantities for other times of the day, and found similar results: high-density grid cells are more packed during daytime than during nighttime, and form larger clusters than during nighttime.
@@ III.3 Critical exponent of cluster size distribution $\tau$ and fractal dimension $D_{f}$
Figure 4(a) shows the probability density function of population cluster size distribution at the critical point, following a power law, for 2:00 P.M. and 2:00 A.M. on April 1st, 2019. We find that the critical exponent $\tau$ = 1.88 at 2:00 P.M. is smaller than that $\tau$ = 2.28 at 2:00 A.M.
@@
A previous study on city growth model using correlated percolation has shown that the exponent of the power-law distribution of urban settlement areas is influenced by long-range correlations between grid cells [32]. Applying this to our results, we can infer that as long-range correlations become stronger, there is a higher probability of larger clusters occurring, resulting in a decrease in the $\tau$ value. Consequently, we can hypothesize that there might be stronger long-range correlation in the spatial distribution of daytime population than nighttime population.
To check the validity of measured critical exponent $\tau$, we employ maximum likelihood fitting methods with goodness-of-fit tests based on the Kolmogorov-Smirnov (KS) statistic and likelihood ratios [33]. When performing a goodness-of-fit test on the real data, we obtain p-values of 0.35 for 2:00 A.M. and 0.10 for 2:00 P.M. These results suggest that it is reasonable to conclude that the distribution follows a power-law behavior in both cases.
The power-law form of cluster size distribution indicates the fractal structure of clusters. Scale-invariant and statistically self-similar clusters are observed in various phenomena from geographical objects such as lakes and islands [34] to urban heat islands [29]. To quantitatively characterize geometrical differences in the clusters at percolation criticality during the daytime and nighttime, we measure the aggregated area - perimeter fractal dimension of the collection of cluster at each occupation probability $p$ using the following equation:
@@
where $n_{p}$ denote the total number of clusters at a given occupation probability $p$, $P_{i}$ and $A_{i}$ are perimeter and area of cluster $i$. The fractal dimension $D_{f}$ typically ranges between 1 and 2. $D_{f}$ approach 2 when clusters have highest degree of irregularity (meaning that the perimeter fills the plane in the cluster) and approach 1 when clusters have smooth shape such as circles [29]. It is also known that, for statistically self-similar surfaces, $D_{f}$ has fractional values between 1 and 2 and these values are independent of occupation probability used for clustering [29, 35].
Figure 5: (a) Aggregated perimeters versus aggregated areas of clusters for occupation probabilities ranging from 0.05 to 0.5, with intervals of 0.05, at 2:00 P.M.(red) and 2:00 A.M.(blue) on April 1st, 2019, demonstrating the same ratio of log(Area) and log(Perimeter). Dashed lines show ratio $D_{f}$ = 1.42 and $D_{f}$ = 1.30 in log - log plot. $D_{f}$ of the perimeter of a circle ($D_{f}$ = 1) and a space-filling plane ($D_{f}$ = 2) are plotted to show the physical bounds for $D_{f}$. (b) Variation of $D_{f}$ April 1st, 2019.
Figure 5(a) shows log - log plots of the aggregated area-perimeter of clusters for multiple occupation probability of $p$ (0.05, 0.1, $\ldots$, 0.5). The data at 2:00 P.M. and 2:00 A.M. exhibit consistent ratios between log(Area) and log(Perimeter length), respectively. We observe that $D_{f}$=$1.42$ at 2:00 P.M. but $D_{f}$=$1.30$ at 2:00 A.M. Figure 5(b) shows the hourly variations in the fractal dimension ($D_{f}$) on April 1st, 2019. Overall, we can observe that $D_{f}$ during daytime is higher than during nighttime.
The reason for these difference of $D_{f}$ is understood by the morphology of large clusters (e.g., the large clusters colored yellow, green, and blue in Fig. 2(e)). As shown in Fig 2(e), these large clusters are more irregular shape and there are more internal voids in the largest cluster(colored yellow) at 2:00 P.M. than at 2:00 A.M.
@@ III.4 Percolation analysis of spatial distribution of population in Helsinki
Figure 6: Results of applying population-based percolation to Helsinki population data (a) Evolution of the largest cluster relative size as a function of $p$ at 2:00 P.M.(red) and 2:00 A.M.(blue). Black lines show the relative size of the largest clusters in the case that occupation is based on the random shuffling of Helsinki population data. (b) Evolution of the number of clusters as a function of $p$ at 2:00 P.M.(red) and 2:00 A.M.(blue). Dashdot lines is number of cluster reaches its maximum, defined as the critical point $p_{c}$. (c) Cluster size distributions at percolation criticality and their power-law scaling exponents at 2:00 P.M.(red circle) and 2:00 A.M.(blue triangle) on workdays. Lines indicate scaling exponent $\tau$ = 1.89 at 2:00 P.M.(red) and $\tau$ = 2.21 at 2:00 A.M.(blue), respectively. (d) Variation of $D_{f}$ during the day. (e) Snapshots of clusters at the critical point for 2:00 P.M. and 2:00 A.M. In the snapshot, we differently colored the top-ranked-size clusters: The yellow cluster represents the first largest cluster. The green cluster represents the second largest cluster. The blue cluster represents the third largest cluster.
@@
First we investigate the size of the largest cluster at 2:00 P.M. and 2:00 A.M. on workdays in Helsinki as a function of occupation probability $p$, respectively as shown in Fig 6(a). As similar to the results obtained from Seoul’s data, the largest clusters in Helsinki start to grow progressively at low occupation probability $p$ and exhibit intermittent jumps. The evolution pattern of the largest cluster size is also different between daytime (2:00 P.M.) and night time (2:00 A.M.). Figure 6(b) show the number of clusters as a function of increasing $p$. The number of clusters increase up to the critical occupation probability $p_{c}$ and decrease thereafter as observed in Seoul.
@@
We also measure the fractal dimension $D_{f}$ of the collection of clusters at criticality for workdays in Helsinki. We observe that $D_{f}=1.43$ at 2:00 P.M. but $D_{f}=1.36$ at 2:00 A.M. In Figure 6(d), the pattern that fractal dimension $D_{f}$ is higher during daytime than during nighttime is consistent with observations from the Seoul data.
The obtained values of $D_{f}$ in Helsinki tend to be higher than in Seoul. This is likely due to the fact that Helsinki has more forests and lakes than Seoul, and it is adjacent to the sea, thus the boundaries of where people can live are more rough in Helsinki than in Seoul.

### v2

# Percolation analysis of spatiotemporal distribution of population in Seoul and Helsinki

@@ Abstract
Spatiotemporal distribution of urban population is crucial to understand the structure and dynamics of cities. Most studies, however, have focused on the microscopic structure of cities such as their few most crowded areas. In this work, we investigate the macroscopic structure of cities such as their clusters of highly populated areas. To do this, we analyze the spatial distribution of urban population and its intraday dynamics in Seoul and Helsinki with a percolation framework. We observe that the growth patterns of the largest clusters in the real and randomly shuffled population data are significantly different, and highly populated areas during the daytime are denser and form larger clusters than highly populated areas during the nighttime. An analysis of the cluster-size distributions at percolation criticality shows that their power-law exponents during the daytime are lower than those during the nighttime, indicating that the spatial distributions of urban population during daytime and nighttime fall into different universality classes. Finally measuring the area-perimeter fractal dimension of the collection of clusters demonstrates that the fractal dimensions during the daytime are higher than those during the nighttime, indicating that the perimeters of clusters during the daytime are rougher than those during the nighttime. Our findings suggest that even the same city can have qualitatively different spatial distributions of population over time, and propose a way to quantitatively compare the macrostructure of cities based on population distribution data.
@@ I Introduction
Characterizing the spatio-temporal distribution of urban population is important because it not only helps us to understand the structure and dynamics of cities but also provides essential information on the demand for traffic, housing, and infrastructure in cities. Spatial distributions of population in many cities are not uniform but quite heterogeneous [1, 2]. A key concept in urban structure, thus, is urban centers or hotspots, usually defined as the most crowded places in cities [3, 4]. Urban structure can be classified as either monocentric or polycentric based on the number of centers and as either compact or dispersed based on how closely clustered centers are [5, 6].
@@
Until recently, the registered population based on censuses has been mainly used to understand the spatial distribution of urban population [1, 7, 8]. These data are highly reliable as collected by the governments, but they are updated every five to ten years and are based on a residence-based population, which limits its ability to describe the dynamics of urban structure. However, real-time or near-real-time population data collected from mobile phone users have recently become available and can be used to better understand urban structure and dynamics [9, 10, 11, 12, 13, 14]. A remarkable work proposed a parameter-free method to identify hotspots in 31 Spanish cities, showed that the number of hotspots in these cities scales sublinearly with their population, and classified these hotspots according to their lifetime [15]. Mobile phone data also allows one to investigate the urban structure encoded in mobility flows of people in cities with the notion of urban centers [16, 17]. These center-based investigations have focused on the microscopic structure of cities although they have revealed important structural properties of cities. A comprehensive understanding of the structure and dynamics of a city requires an investigation of the macrostructure of the spatial distribution of the city’s population such as clusters of highly populated areas.
@@
Percolation can provide an effective way to investigate the macroscopic structure underlying a given spatial distribution of variables of interest from temperature and altitude to congestion and population [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28]. For example, a work that analyzed spatial distribution of temperature in cities with a temperature-based percolation (i.e., occupying sites in order of the highest temperature to the lowest temperature) revealed that urban heat islands are fractal despite the diversity in urban form, composition, and environment across the world [29]. Another work, using a percolation model to analyze spatial distributions of congestion in two cities, showed that these distributions belong to different universality classes of percolation depending on the commuting patterns of the cities [20]. Although many previous works on urban structure have focused on center-based approaches, these have limitations in revealing urban macrostructure[9, 30, 31, 32, 33]. With a percolation-based framework, we can not only understand urban macrostructure in terms of clusters but can also quantitatively compare macroscopic characteristics of urban structure, for example, using critical exponents and fractal dimensions.
@@
In this work, we investigate the intraday dynamics of the macroscopic structure underlying the spatial distribution of populations in Seoul and Helsinki with a percolation framework. We examine the characteristics of clusters of highly populated areas and the connectivity between these clusters as they are created by occupying areas from high to low population in this framework. We observe that the growth patterns of the largest clusters in the real-world and randomly shuffled population data are significantly different, and highly populated areas during the daytime are denser and form larger clusters than highly populated areas during the nighttime by analyzing the size of the largest clusters, the radius of gyration of occupied areas, and the correlation length of clusters in the data. We investigate the cluster-size distributions at the percolation critical point and observe that their power-law exponents during daytime are lower than during nighttime, indicating that the spatial distributions of urban population during daytime and nighttime belong to different universality classes. We measure the area-perimeter fractal dimension of the collection of clusters and observe that the fractal dimensions during the daytime are higher than those during the nighttime, indicating that the perimeters of high-population clusters during daytime are rougher than during nighttime.
@@ II.1 Data
We use the Seoul Living Population Data (SLPD) [34] provided by the Seoul Metropolitan Government to analyze the spatio-temporal distribution of population in Seoul. The SLPD is based on mobile phone signal data and tells us the total number of people of a specific area in Seoul at a specific time (an hourly interval). There are 19,153 area units, called ‘jibgyegu’, in the dataset. Figure 1(a) shows the spatial distribution of Seoul’s population at 9:00 A.M. on April 1, 2019 by the area unit.
Since the ‘jibgyegu’ area units are irregular in shape and size, we spatially aggregated the number of people recorded each hour in each area unit on a regular square grid of cell size 400 m. As a result, we obtained 4043 grid cells with geographic coordinates and hourly population density. We chose a grid cell size of 400 m to get a sufficient number of grid cells for statistical analysis but also to minimize the impact of noises from too-small size and the impact of natural boundaries such as mountains and rivers in Seoul. We also tested the robustness of our results with two different sizes of grid cells (250 and 500 m) and obtained qualitatively similar results. We considered weekdays in April 2019 for our analysis (20 days in total).
To focus on relatively high population areas in cities, we excluded grid cells whose daily population densities are lower than the daily population densities averaged over all the grid cells. Figure 1(b) shows the spatial distribution of Seoul’s population at 9:00 A.M. on April 1, 2019 on the grid cells after the exclusion of lowly-populated cells. The number of grid cells was reduced from 4043 to 2480. The excluded lowly-populated grid cells are mostly located in mountains, rivers, and public facilities such as sewage treatment plants.
Figure 1: Spatial distribution of population density in Seoul at 9:00 A.M. on April 1, 2019. The color bar represents the logarithmic values of population density in spatial units. (a) Spatial distribution by the ‘jibgyegu’ units. (b) Spatial distribution by grid cells of 400 m size after the exclusion of lowly-populated cells.
@@ II.2 Population-based percolation
The procedure of our population-based percolation is as follows. We first rank all the grid cells according to their population density from the highest to the lowest value. Next, we occupy the grid cells one by one according to their rank and observe how clusters evolve with the occupation probability $p$. A cluster, defined as a set of nearest neighboring occupied grid cells, represents a densely populated macroscopic area in a city. The occupation probability $p$ is defined as the number of occupied grid cells divided by the total number of grid cells.
@@ III.1 Evolution of the largest cluster and the number of clusters
Our percolation analysis aims at characterizing the macroscopic structure underlying spatial distribution of urban population by investigating the clusters of highly-populated areas in cities. First we investigate the size of the largest cluster as a function of the occupation probability $p$. Figure 2(a) and 2(b) show the size of the largest cluster at 2:00 P.M. and 2:00 A.M. on April 1, 2019 in Seoul as a function of $p$, respectively.
@@ III.2 Radius of gyration and correlation length
We considered these quantities for other times of the day, and found similar results: high-density grid cells during the daytime are more packed and form larger clusters than high-density grid cells during the nighttime.
@@ III.3 Critical exponent of cluster size distribution $\tau$ and fractal dimension $D_{f}$
Figure 4(a) shows the probability density function of population cluster size distribution at the critical point, following a power law, for 2:00 P.M. and 2:00 A.M. on April 1st, 2019. We find that the critical exponent $\tau$ = 1.89 at 2:00 P.M. is smaller than that the critical exponent $\tau$ = 2.28 at 2:00 A.M.
@@
A previous study on a city growth model using correlated percolation has shown that the exponent of the power-law distribution of urban settlement areas is influenced by long-range correlations between grid cells [36]. Applying this to our results, we can infer that as long-range correlations become stronger, there is a higher probability of larger clusters occurring, resulting in a decrease in the $\tau$ value. Consequently, we can hypothesize that there might be a stronger long-range correlation in the spatial distribution of daytime population than in that of nighttime population.
To check the validity of the measured critical exponent $\tau$, we employ maximum likelihood fitting methods with goodness-of-fit tests based on the Kolmogorov-Smirnov (KS) statistic and likelihood ratios [37]. When performing a goodness-of-fit test on the real data, we obtain p-values of 0.35 for 2:00 A.M. and 0.10 for 2:00 P.M. These results suggest that it is reasonable to conclude that the distribution follows a power-law behavior in both cases.
The power-law form of cluster size distribution indicates the fractal structure of clusters. Scale-invariant and statistically self-similar clusters are observed in various phenomena from geographical objects such as lakes and islands [38] to urban heat islands [29]. To quantitatively characterize geometrical differences in the clusters during the daytime and nighttime, we measure the aggregated area - perimeter fractal dimension of the collection of cluster at each occupation probability $p$ using the following equation:
@@
where $n_{p}$ denotes the total number of clusters at a given occupation probability $p$, and $P_{i}$ and $A_{i}$ are the perimeter and the area of cluster $i$. We define the perimeter of a cluster as the whole boundary between the cluster and the adjacent nonoccupied grid cells. The fractal dimension $D_{f}$ typically ranges between 1 and 2. $D_{f}$ approach 2 when clusters have the highest degree of irregularity (meaning that the perimeter fills the plane in the cluster) and approach 1 when clusters have a smooth shape such as a circle [29]. It is also known that, for statistically self-similar surfaces, $D_{f}$ has fractional values between 1 and 2 and these values are independent of the occupation probability used for clustering [29, 39].
Figure 5: (a) Aggregated perimeters versus aggregated areas of clusters for occupation probabilities ranging from 0.05 to 0.5, with intervals of 0.05, at 2:00 P.M.(red) and 2:00 A.M.(blue) on April 1st, 2019, demonstrating the same ratio of log(Area) and log(Perimeter). Dashed lines show ratio $D_{f}$ = 1.42 and $D_{f}$ = 1.30 in the log - log plot. $D_{f}$ values of the perimeter of a circle ($D_{f}$ = 1) and a space-filling plane ($D_{f}$ = 2) are plotted to show the physical bounds for $D_{f}$. (b) Variation of $D_{f}$ April 1st, 2019.
Figure 5(a) shows log - log plots of the aggregated area-perimeter of clusters for multiple occupation probability of $p$ (0.05, 0.1, $\ldots$, 0.5). The data at 2:00 P.M. and 2:00 A.M. exhibit consistent ratios between log(Area) and log(Perimeter length), respectively. We observe that $D_{f}$=$1.42$ at 2:00 P.M. but $D_{f}$=$1.30$ at 2:00 A.M. Figure 5(b) shows the hourly variations in the fractal dimension ($D_{f}$) on April 1, 2019. Overall, we can observe that $D_{f}$ during the daytime is higher than that during the nighttime.
The reason for these differences of $D_{f}$ is understood by the morphology of large clusters (e.g., the large clusters colored yellow, green, and blue in Fig. 2(e)). As shown in Fig. 2(e), these large clusters are more irregular in shape and there are more internal voids in the largest cluster(colored yellow) at 2:00 P.M. than at 2:00 A.M.
@@ III.4 Percolation analysis of spatial distribution of population in Helsinki
Figure 6: Results of applying population-based percolation to Helsinki population data (a) Evolution of the largest cluster relative size as a function of $p$ at 2:00 P.M.(red) and 2:00 A.M.(blue). Black lines show the relative size of the largest clusters in the case that occupation is based on the random shuffling of Helsinki population data. (b) Evolution of the number of clusters as a function of $p$ at 2:00 P.M.(red) and 2:00 A.M.(blue). Dashdot lines is number of cluster reaches its maximum, defined as the critical point $p_{c}$. (c) Cluster size distributions at percolation criticality and their power-law scaling exponents at 2:00 P.M.(red circle) and 2:00 A.M.(blue triangle) on workdays. Lines indicate scaling exponent $\tau$ = 1.88 at 2:00 P.M.(red) and $\tau$ = 2.21 at 2:00 A.M.(blue), respectively. (d) Variation of $D_{f}$ during the day. (e) Snapshots of clusters at the critical point for 2:00 P.M. and 2:00 A.M. In the snapshot, we differently colored the top-ranked-size clusters: The yellow cluster represents the first largest cluster. The green cluster represents the second largest cluster. The blue cluster represents the third largest cluster.
@@
First we investigate the size of the largest cluster at 2:00 P.M. and 2:00 A.M. on workdays in Helsinki as a function of occupation probability $p$, respectively as shown in Fig 6(a). Similar to the results obtained from Seoul’s data, the largest clusters in Helsinki start to grow progressively at low occupation probability $p$ and exhibit intermittent jumps. The evolution pattern of the largest cluster size is also different between daytime (2:00 P.M.) and night time (2:00 A.M.). Figure 6(b) show the number of clusters as a function of increasing $p$. The number of clusters increase up to the critical occupation probability $p_{c}$ and decrease thereafter as observed in Seoul.
@@
We also measure the fractal dimension $D_{f}$ of the collection of clusters for workdays in Helsinki. We observe that $D_{f}=1.43$ at 2:00 P.M. but $D_{f}=1.36$ at 2:00 A.M. In Figure 6(d), the pattern that fractal dimension $D_{f}$ is higher during daytime than during nighttime is consistent with observations from the Seoul data.
The obtained values of $D_{f}$ in Helsinki tend to be higher than in Seoul. This is likely due to the fact that Helsinki has more forests and lakes than Seoul, and it is adjacent to the sea, thus the boundaries of where people can live are more rough in Helsinki than in Seoul. The critical exponent $\tau$ and fractal dimension $D_{f}$ of Seoul on April 1, 2019 and Helsinki on workdays are summarized in Table I for comparison.
Critical exponent $\tau$ Area-perimeter fractal dimension
Seoul (2:00 P.M.) 1.89 1.42
Seoul (2:00 A.M.) 2.28 1.30
Helsinki (2:00 P.M.) 1.88 1.43
Helsinki (2:00 A.M.) 2.21 1.36
Table 1: Critical exponent $\tau$ and fractal dimension $D_{f}$ for the considered cities and time.